Macroeconomics I need help please with b and c

media%2F2eb%2F2eb438e2-f7a8-4ead-8276-e5
Answer #1

(b)

If we move from point C (toward point D),

Cost of one more automobile = Change in number of forklifts / Change in number of automobiles = (21 - 12) / (6 - 4) = 9/2 = 4.5

Cost of one more forklift = Change in number of automobiles / Change in number of forklifts = (6 - 4) / (21 - 12) = 2/9 = 0.22

Law of increasing opportunity cost is reflected in the shape of PPC. If the PPC is bowed out in shape, there is  increasing opportunity cost. If the PPC is linear, there is constant opportunity cost.

(c)

When the economy produces 2 automobiles, it produces 27 forklifts and when the economy produces 4 automobiles, it produces 21 forklifts. Therefore, when the economy produces 3 automobiles, it produces less than 27 but more than 21 forklifts. If it produces 20 forklifts for 3 automobiles being produces, the economy is under-producing and not utilizing all the resources. Such a point will lie inside the PPC, reflecting inefficiency in production.